OpenStudy (anonymous):

Write an equation of a parabola with a vertex at the origin and a focus at (–2, 0). Answer

OpenStudy (anonymous):

x = a(y - k) + k a = 1/(4p) a = -1/8 x = -1/8y^2 This is because there is no phase shift or vertical chift